OpenStudy (anonymous):

Molly has x marbles. If she adds y marbles, she has a total of 14. However, if she took her x number of marbles and subtracted y marbles, she only has 8. What is the value of x and y? Write your answer as an ordered pair. (11, 3) (10, 2) (8, 14) (3, 11)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no no no it's probably A

OpenStudy (dbzfan836):

you sure? check your work

OpenStudy (anonymous):

for clarity, (not knowing what class this is), I would set the two equations equal to each other: \[x+y=14\] subtract 14 from both sides \[x+y-14=0\] \[x-y=8\] and then subtract 8 from both sides \[x-y-8=0\] then combine \[x+y-14=x-y-8\] solve \[x+y+[-14+14]=x-y+[14-8]\]

OpenStudy (anonymous):

\[x+y=x-y+6\] \[[x-x]+[y+y]=[x-x]+[-y+y]+6\] \[2y=6\] \[\frac{ 2 y}{ 2 }=\frac{ 6 }{ 2 }\] \[y=3\] plug that back into the original \[x+3=14\] \[x=11\]

OpenStudy (anonymous):

Thank you